# Fine-grained Preference Optimization Improves Zero-shot Text-to-Speech

### v1

@@ 5.2 Subjective Evaluation
Apart from NMOS, we also conduct ABX preference tests to verify the performance gains of our proposed FPO compared to baseline systems. As shown in Figure 4, compared to the base model, FPO is preferred in 54.3% of cases and ties in 24.1%, significantly outperforming the base model in preference. Notably, FPO also surpasses SDPO by a considerable margin (44.2% vs. 28.7%, with 27.1% ties), demonstrating the effectiveness of fine-grained optimization compared to full-sequence optimization. The ABX test results further highlight FPO’s advantages.
@@
These results indicate that while utterance-level optimization methods effectively align the base model to some extent, FPO directly addresses critical pain points in the human listening experience. This leads to higher practical utility, even when objective metric scores are comparable. The strong alignment between subjective and objective evaluations further validates the effectiveness of the proposed FPO.
5.3 Analysis on Data Efficient
@@
To further evaluate the data efficiency of our proposed FPO, we compare its performance with baseline systems using small-scale Mandarin training data, with CER and bad case ratio as evaluation metrics. As shown in Figure 5, FPO achieves a similar CER performance as the RIO system while using only 200 utterances, compared to the 800 utterances required by the baselines. FPO is up to four times faster than the baseline systems. Additionally, FPO outperforms all baselines in bad case ratio across different data sizes and achieves a speedup of approximately 4x compared to UNO.
Furthermore, we scale up the preference data size to compare the data scalability between the SDPO approach and our proposed FPO. As shown in Table 2, with just 1,000 training utterances, FPO reduces CER by 52.5% (from 8.26 to 3.92) and WER by 54.8% (from 6.97 to 3.15) for CN and EN, respectively. This outperforms even SDPO trained on 5,000 utterances (CER: 4.14 vs. 3.11; WER: 3.85 vs. 2.56), highlighting FPO’s ability to maximize efficiency with limited data. Notably, FPO reduces the bad case ratio to 8% (CN) and 5% (EN) with only 1,000 samples, whereas SDPO requires 5,000 samples to achieve 9% (CN) and 8% (EN). These results further demonstrate FPO’s effectiveness in mitigating critical errors while maintaining high data efficiency.
@@
Table 2: Comparison results between FPO and SDPO on several larger datasets. "Num" represents the number of utterances used for training.
Num (#) CER/WER Bad case ratio
CN EN CN EN
CosyVoice - 8.26 6.97 31% 24%
SDPO 1000 7.76 6.54 17% 13%
2000 6.06 5.79 14% 12%
5000 4.14 3.85 9% 8%
FPO 1000 3.92 3.15 8% 5%
2000 3.79 2.84 5% 4%
5000 3.11 2.56 3% 2%
FPO’s advantage stems from its selective training loss strategy, which focuses optimization on error segments identified through fine-grained annotations. By avoiding unnecessary updates to well-learned tokens, FPO reduces computational redundancy and accelerates convergence. In contrast, utterance-level methods like SDPO apply uniform optimization across entire sequences, diluting the impact of critical segments and requiring larger datasets to achieve comparable improvements.
6 Conclusion
In this study, we propose FPO, a novel RLHF framework designed to address localized errors in zero-shot TTS systems. By categorizing generation errors into systemic and segmental types and introducing token-level selective loss optimization, FPO aligns problematic segments using fine-grained preference data while avoiding computational waste on well-learned tokens. Experimental results demonstrate that FPO significantly improves speech intelligibility and robustness, reducing CER by 52.5% and WER by 54.8% compared to the base model. Additionally, FPO lowers the bad case ratio to 8% (CN) and 5% (EN), outperforming utterance-level preference optimization approaches. Subjective evaluations further confirm FPO’s superiority in naturalness and alignment with human perceptual priorities. Notably, FPO achieves comparable performance with 4× fewer training samples than baseline systems, highlighting its remarkable data efficiency.
Limitations
@@ 6 Conclusion
One of the main limitations of this study lies in the construction of the preference training dataset. The proposed framework relies on fine-grained manual annotation to ensure the quality and consistency. This process is both time-consuming and labor-intensive, which may limit the scalability of the approach to larger or more diverse datasets. Although our proposed approach is data efficient and need less annotation data compared with previous works, the fine-grained annotation at each sample is more difficult than utterance level annotation. In future work, we plan to explore strategies to alleviate this limitation. For example, leveraging weakly supervised, semi-supervised, or self-supervised learning techniques may help reduce the dependence on manually annotated data.
Ethics Statement
This work does not raise any ethical concerns. All data and pre-trained models used in this study are publicly available and comply with the following licenses: Creative Commons BY 4.0, Creative Commons CC0, Creative Commons BY-NC-ND 4.0, and Creative Commons BY-SA 4.0.
Appendix A Appendix
A.1 Segmental Errors
Temporal modeling errors stem from the model’s limitations in capturing the temporal dynamics of speech, leading to issues such as mispronunciation, where abrupt or unstable pronunciation changes occur due to insufficient modeling of pitch continuity or variation trends. Another common problem is abnormal silence, where the model erroneously inserts silent frames or extends silent segments, disrupting the naturalness of generated speech. Additionally, the generated speech may exhibit unnatural prosody (i.g. unnatural pause), as the model may fail to adequately capture local rhythmic patterns, including stress, duration, and speaking rate, resulting in speech that deviates from natural prosody.
Semantic-phonetic alignment errors arise from mapping between the phonetic input and the acoustic output. These errors often manifest as repetition, where the model over-relies on its previous outputs and redundantly generates the same words or phonemes. Alternatively, truncation errors may occur when the language model fails to encode key phonetic information from the input or neglects certain critical elements, leading to synthesized speech that lacks completeness and accuracy.
A.2 Prosody Annotation
For prosody issues, we primarily focus on abnormal pauses and annotate fine-grained prosody errors through a three-step process. First, we use an automatic speech recognition system to transcribe the speech into text. Next, we employ text LLMs to determine ideal pauses in the text and insert break tags. Finally, professional annotators evaluate the speech samples based on the break tags, identifying preferred and dispreferred prosody samples. Fine-grained timestamps are then obtained using force-alignment tools based on the text and break tags.

### v2

@@ V-B Subjective Evaluation
In addition to NMOS, we conduct ABX preference tests to further validate the performance gains of our proposed FPO compared to baseline systems. The ABX tests include both CosyVoice and CosyVoice2 models, while all other baselines are trained based on CosyVoice2. As shown in Figure 3, FPO is preferred over CosyVoice in 57.1% of cases and ties in 21.4%. Compared to CosyVoice2, FPO achieves a 44.3% win rate, again significantly outperforming the base model in terms of listener preference. Notably, FPO also outperform SDPO by a substantial margin (40.8% vs. 27.6%, with 31.6% ties), demonstrating the effectiveness of fine-grained optimization over full-sequence optimization. These ABX test results further emphasize the advantages of FPO in producing more preferred and natural speech.
@@
These results indicate that while utterance-level optimization methods effectively align the base model to some extent, FPO directly addresses critical points in the human listening experience. This leads to higher practical utility, even when objective metric scores are comparable. The consistency between subjective and objective evaluations further validates the effectiveness of the proposed FPO.
V-C Analysis on Data Efficient
@@
To further evaluate the data efficiency of our proposed FPO, we compare its performance against baseline systems using small-scale English training data, with intelligibility evaluated on the test-hard set using WER as metric and average bad case ratio across three test sets as another metric. As shown in Figure 4, FPO achieves comparable WER performance to the UNO system using only 200 training utterances, whereas the baselines require 600 utterances. This demonstrates that FPO is up to three times more data-efficient than baseline systems. Furthermore, FPO consistently outperforms all baselines in terms of bad case ratio across various data scales and achieves an approximate 4x speedup compared to UNO.
@@
TABLE III: Comparison results between FPO and SDPO on several larger-scale datasets. ”Num” represents the number of utterances used for training.
Num (#) CER WER Bad case ratio
CosyVoice2 - 1.45 5.33 14%
 + SDPO 1000 1.43 5.52 15%
2000 1.40 5.19 12%
5000 1.31 3.94 7%
 + FPO 1000 1.32 4.58 8%
2000 1.29 3.76 5%
5000 1.21 3.23 3%
Furthermore, we scale up the preference data size to compare the data scalability between the SDPO approach and our proposed FPO. As shown in Table III, with just 1,000 training utterances, FPO reduces the CER of the original CosyVoice2 model from 1.45 to 1.32 and the WER from 5.33 to 4.58 for Mandarin and English evaluation datasets, respectively. This performance outperforms even SDPO trained on 5,000 utterances (CER: 1.32 vs. 1.31; WER: 4.58 vs. 3.94), highlighting FPO’s ability to maximize efficiency of preference optimization with limited data. Notably, FPO reduces the bad case ratio from 14% to 8% using only 1,000 samples, whereas SDPO requires 5,000 samples to achieve a similar performance of about 7%. These results further demonstrate FPO’s effectiveness in mitigating critical errors while maintaining high data efficiency.
FPO’s advantage stems from its selective training loss strategy, which precisely directs optimization toward error-prone segments identified through fine-grained annotations of temporal modeling and semantic-phonetic misalignments. By identifying problematic tokens and excluding well-learned regions from gradient updates, FPO minimizes computational redundancy while preventing over-optimization of already proficient segments. This targeted approach not only reduces noise in parameter updates but also accelerates convergence by prioritizing problematic segments. In contrast, utterance-level methods like SDPO apply uniform optimization across entire sequences, inadvertently diluting the impact of critical segments by averaging gradients from both optimal and suboptimal regions. This indiscriminate updating forces such methods to rely on larger datasets (4–5× more samples in ablation studies) to statistically filter irrelevant signals and achieve comparable error reduction. Furthermore, the utterance-level optimization strategy risks destabilizing well-trained segments through unnecessary perturbations, whereas FPO’s surgical precision preserves existing capabilities while efficiently resolving localized failures.
TABLE IV: Comparison between our proposed FPO and other preference alignment approaches.
SDPO UNO RIO Mask-DPO FPO
Grain Utterance Utterance Utterance Sentence Token
Goal Overall quality Uncertainty perceptual Robustness enhancement Factuality alignment Targeted correction
Annotation Global preference Global preference Global preference Global preference Token-level preference
Reward Binary Binary Binary Binary Fine-grained
VI Limitations and Discussion
@@
One of the main limitations of this study lies in the construction of the preference training dataset. The proposed framework relies on fine-grained manual annotation to ensure the quality and consistency. This process is both time-consuming and labor-intensive, which may limit the scalability of the approach to larger or more diverse datasets. Although our proposed approach is data efficient and need less annotation data compared with previous works, the fine-grained annotation at each sample is more difficult than utterance level annotation. Furthermore, our proposed approach is not inherently restricted to autoregressive TTS models. Beyond the evaluated autoregressive backbones, the framework can be naturally extended to discrete non-autoregressive models. In these cases, the underlying mechanism of preference-based optimization remains directly applicable without substantial modification. For TTS paradigms that rely on continuous representations, the core idea of our method is still conceptually applicable. However, applying the method in this setting introduces additional challenges. In particular, constructing sufficiently diverse and well-controlled paired samples from continuous representations for effective reinforcement learning becomes non-trivial. Meanwhile, practical challenges may arise due to differences in training objectives and generation mechanisms, which we leave for future work.
VII Conclusion
In this study, we propose FPO, a novel RL framework that introduces a new paradigm for enhancing zero-shot TTS robustness through localized error correction. By systematically categorizing generation errors into temporal modeling discrepancies and semantic-phonetic misalignments, and applying token-level preference optimization, FPO adopts a segment-focused strategy that significantly outperforms baseline systems. This validates our hypothesis that targeted intervention on error-prone segments leads to more efficient and effective error correction. FPO reduces the bad case ratio by 12% compared to base models through fine-grained perceptual alignment. Additionally, it achieves comparable performance using fewer training samples than baseline systems, demonstrating strong data efficiency. Experimental results confirm that FPO substantially improves the robustness of the base TTS system, while subjective evaluations further support FPO’s superiority in naturalness and alignment with human perceptual preferences.